sequences and series question from special series
find the sum to n terms n2+2 
 

Grade:11

1 Answers

BALAJI ANDALAMALA
8 years ago
\sum_{n=1}^{n}(n^2+2) = \sum_{n=1}^{n}n^2+\sum_{n=1}^{n}2
= \frac{n(n+1)(2n+1)}{6}+2n
= \frac{n(2n^2+3n+13)}{6}